# Beyond the Main Sequence: Binary Evolution Pathways to Blue Stragglers in the Gaia Era

### v1

@@ 3.1 Absolute Numbers and Fractions of BSSs
Figure 3 shows the absolute number () of detected BSSs and their number ratio relative to MS stars () in each of our sample GCs and OCs as a function of the parent cluster age. For the denominator of , we use the number of MS stars in the luminosity range between the MSTO and MSTO+1 mag, similar to previous studies (JS21, Cordoni et al. 2023). We select MS stars using the same criteria for all clusters shown in the bottom panel of Figure 3 to guarantee a homogeneous comparison. Due to the larger average distances of GCs and, therefore, their fainter MSTOs, our earlier photometric depth cut (see Sect. 2.2) collides with the fainter bound of the luminosity range definition for some distant GCs. For this reason we have inspected by eye every CMD in our GC sample and only provide values for those clusters whose MS stars at fainter than the MSTO are within the photometric depth bounds. With this limitation, we are able to obtain values for 21 of 41 sample GCs, which is the reason why the number of GCs in the bottom panel of Figure 3 is smaller than the corresponding number in the upper panel. We have also listed the values of , and in Table 2 along with other stellar parameters of the clusters.
@@
The inset panels show zoom-ins at the oldest ages to better visualize any localized trends for the GC sub-sample. In Figure 3, we also compare the number of BSSs identified in this work against other Gaia-based studies from the literature, i.e. R21, JS21, Li et al. (2023) and Cordoni et al. (2023). We point out that our sample is the first that homogeneously studies BSSs in GCs and OCs. We observe that in all studies, including ours, the BSS population in OCs tends to increase in absolute and fractional numbers towards older cluster ages. However, this trend appears to reach a maximum at GC ages (), and might even be consistent with the numbers of identified BSSs decreasing with increasing age slightly. We also perform sliding histograms of the distributions of () and (dashed line) in Figure 3. shows a clear trend of increasing with age, while the distribution appeared as a plateau in the range before suddenly rising again with the varying BSS fraction at older ages. We also perform the Pearson’s r-coefficients in the Monte Carlo method 111 Pearson’s test measures whether a pair of two variables are linearly correlated or not. The r-coefficient is defined as the covariance of the variables divided by the product of their standard deviations, and the value ranges from 1 to 1. 1 signifies perfect positive linear correlation between the variables, 1 signifies perfect linear anti-correlation, and 0 means no linear correlation. Based on Cohen (1988), coefficient values 0.5 (0.5) indicate a strong positive (negative) linear relation, values between 0.3 and 0.5 (0.3 and 0.5) indicate a moderate positive (negative) linear relation, and values between 0 and 0.3 (0 and 0.3) indicate a weak positive (negative) linear relation. The Pearson’s test also provides a p-value that indicates whether the observed correlation is statistically significant or not. It depends on the number of data points or the sample size. A smaller p-value (generally 0.05) suggests that the correlation is statistically significant, while for larger values (0.05), we can not conclude a significant correlation (Wasserstein & Lazar, 2016). (to include observational errors) for both distributions (Curran, 2014; Privon et al., 2020)222https://github.com/privong/pymccorrelation?tab=readme-ov-file and find that the values are 0.67 for and 0.45 for .333The numbers in all cases denote the correlation coefficient 50th percentile. The lower and upper errors denote the 16th and 75th percentiles, respectively. The p-values for both cases are found to be 0. The coefficients and p-values suggest that and , both correlate with the age of the clusters. Although the correlation for is moderate, one cannot deny the sudden rise of BSS fraction at . The plateau in in the range could be because, at a young age, massive stars reach larger radii in the AGB phase and mass transfer is possible even in wide binaries, which favours BSS formation to make a constant BSS fraction with respect to the MS population (Leiner & Geller, 2021). Other possible mechanisms of more BSS formation are wind accretion from the AGB companion and stable mass transfer in the subgiant branch or Hertzprung gap in younger clusters. However, the increase in BSS fractions at older ages could be due to the longer MS lifetime of low-mass BSSs in older clusters and/or the contribution from collisional BSSs (Leiner & Geller, 2021, see their Fig.5).
@@
However, Pearson’s r-coefficient () suggests that the anti-correlation is rather weak. A similar weak anti-correlation () is also found for as well. The p-values are found to be 0 for both cases. As the very central regions of the GCs are excluded in this study because of Gaia’s limitation, incorporating HST observations alongside Gaia data would be beneficial for investigating this correlation. At the same time, we need to include more number of GCs with deep photometry, as we missed to include some GCs due to the magnitude limitation of Gaia.
@@
3.2 BSS Astrophysical Parameters
We derive the following astrophysical parameters for our sample BSSs: effective temperature (), stellar mass (), luminosity () and surface gravity (). We use three different approaches: the infrared flux method (IRFM; from color-temperature relations), the isochrone best–fit method (from PARSEC isochrones) and the XPP from Gaia DR3 spectra, if available. While the isochrone method provides all astrophysical parameters, the IRFM method provides and the XPP method gives & . We mainly use the parameters obtained from the isochrone method however, it is necessary to compare values, obtained from all three methods to check the consistency in the parameter estimations. In Appendix B, we describe all three methods and the comparison of the values obtained from them, in detail. In the following subsections, we summarize the results and compare the stellar parameters with the literature values.
@@ 3.2.1 Effective Temperature
We notice that values extracted from different approaches range from 6000 K to 9000 K and appear consistent with each other with a little systematic offset towards higher temperature ( ). From these comparisons, we estimate that the systematic uncertainty between various methods is of the order K. From the IRFM we find for GCs and for OCs. From isochrone fitting we find for GCs and for OCs. Only 537 out of 4399 BSSs ( of our sample) have XPP parameters measured for 279 stars in GCs and 258 stars in OCs. With this method, we find the mean BSS effective temperature of for GCs and for OCs. We have listed these parameters in Table 3 and the full table will be available online in an electronic format.
Table 1 compare our derived values for with those found in the literature for various OCs and GC, as well as BSS samples. We find that our effective temperatures derived are in very good agreement with previous studies. For example, Jadhav et al. (2019), Sindhu et al. (2019) and Pandey et al. (2021) have analyzed different BSSs in the OC NGC 2682 (also known as M67) by fitting models to spectral energy distributions (SEDs) searching for BSS hot companions; these studies derive BSS temperatures in the range between and , which are in excellent agreement with our measurements, yielding a mean for all methods combined. We also point out the excellent agreement with the values derived in Gosnell et al. (2015), where the authors find for 19 BSSs, which is in very good agreement with our results independent of the applied method ( and , for the IRFM and isochrone method, respectively). Finally, our is also in very good agreement with theoretical predictions for BSSs in GCs from Stȩpień & Kiraga (2015), who calculate , and who had applied different initial conditions to binary systems.
@@ 3.2.2 Stellar Mass and Age of the BSSs
Stellar mass and age are important parameters in the study of the evolutionary state of BSSs. We derive these parameters along with equivalent stellar luminosity, and stellar surface gravity for all BSSs in our sample based on the isochrone fitting method. We obtain the mean luminosity for GCs and for OCs, and find the mean surface gravity of for GCs and for OCs. These values correspond to a mean stellar mass for BSSs in GCs and for BSSs in OCs, which altogether translates to a mean stellar age of for BSSs in GCs and for BSSs in OCs.
We compare these values with previous results from the literature in Table 1. We find that the average mass of our BSS sample in GCs () differs from previous works, which report (Shara et al., 1997) and (Baldwin et al., 2016). This may be partly explained by the lower BSS numbers of previous studies, but also by either of the previous arguments, i.e. i) limitations of the Gaia data selection and ii) BSS mass segregation in GCs. In particular, the latter argument lends credence to the reason for this offset as the two previous studies are based on HST observations which target the GC centres. Our sample, on the other hand, is based on BSS samples from outside the half-light radius. Nevertheless, as can be appreciated in Table 1, our BSS mass results are in very good agreement with those measured by Simunovic (2017, ), Fiorentino et al. (2014, ) and Stȩpień & Kiraga (2015, ). The comparison with Simunovic (2017) is in particular striking, as this work measured the parameters for 1507 BSSs in 38 GCs using the isochrone fitting technique similar to our procedure, which makes it statistically the most robust comparison.
@@
Notes. For data provided by Jadhav & Subramaniam (2021, or JS21) we consider stars labeled as “‘BSS” and “‘pBSS” (probable BSS).
@@
The average BSS mass for our OC sample ( compares favourably with the corresponding value found by JS21 considering the uncertainties. They report an average-mass value of if only confirmed BSSs from their catalog are considered, and if both confirmed and candidate/probable BSSs listed in their catalog are included. As was explained previously, the difference between these works is likely the result of our Gaia magnitude cut-off criterion, discarding all bright stars with as they are likely saturated and, therefore, their Gaia parameters likely affected by systematic errors. This not only explains the difference between our results, as well as why our dispersion in BSS mass for OCs is smaller, but also the lack of BSSs younger than Myr in our sample, most of which would naturally be among the most massive BSSs.
3.3 Effective temperature vs cluster age
@@
In Figure 4 we compare the average MSTO offset in of BSS in a cluster against the corresponding cluster age. To compare self-consistently all the BSSs with respect to the same evolutionary reference point, given the clusters’ very different ages and metallicities, we use the of their MSTO as reference. For each BSS, we obtain . We then compute which is the mean difference between the BSS temperatures and the in each cluster. As XPP parameters are not available for stars at the MSTO, we only do this for obtained by isochrones and color–temperature relation, i.e. IRFM, and recall the systematic offsets found in the comparison of these methods (see Fig. 14). The errorbars in represent the dispersion of BSS temperatures in each cluster, while the errorbars on the horizontal axis are age uncertainties.
@@
In the top panel of figure 4, BSSs in OCs show on average a trend against the parent cluster age where younger OCs show larger offsets, i.e. hotter BSSs, with respect to those found in older OCs, or in other words, the difference between the effective temperature of BSSs and the MSTO effective temperature decreases with increasing parent cluster age. Pearson’s coefficients for IRFM (isochrone) method is (), suggesting the correlation to be weak, which could be due to a large dispersion in the BSS temperatures in the younger cluster. However, we expect the observed trend since younger clusters host more massive stars which are also larger in diameter (e.g. Iben, 2013), and therefore, have likely more efficient MT between binary members and higher collision/capture probabilities with other stars. These hotter BSS stars evolve faster than BSS stars around the MSTO and, thus, the difference in average between BSSs and BSSs around MSTO is expected to decrease with time. We also notice that the dispersion in decreases with the age of the OC, which may indicate that BSSs tend to be more concentrated in the color-luminosity space as the OC ages, which is also shown in the bottom panel of the figure. The Pearson’s coefficient values further confirm the trend: and for IRFM and isochrone methods, respectively.
@@
We do not observe any of these correlations for BSSs in GCs, which is also supported by the Pearson’s coefficient for IRFM (isochrone) method () (see inset panels in Fig. 4). However, we must keep in mind that we might be exposed to biases for the GC sub-sample: i) we have applied a color cut when selecting BSSs in GCs to avoid including BHB stars (see Fig. 2 and Sect. 2.2) which, although unlikely may cut into the distribution of BSSs with extremely large (this could in principle be tested with UV photometry); ii) Gaia is only capable to observe stars outside (half-light radius) in GCs, which could change the mean if we assume that mass segregation has already acted on the BSS population (Contreras Ramos et al., 2012; Ferraro et al., 2012). Without consideration of these potential biases and taking the distributions in Figure 4 at face value, this means that as GCs age past Gyr, BSSs are more or less equally distributed in .
@@
In the middle panel of figure 4, we check if fractional excess with respect to has any correlation with the cluster age. We do not find any trend for either OCs or GCs, which is also supported by the Pearson’s coefficient test. Pearson’s coefficients for IRFM (isochrone) are () for OCs and ().
3.4 BSS Fractional Mass Excess
@@
BSSs are expected to gain mass during their MS lifetime. This means that the “excess” between their current mass and the stellar mass at the MSTO could provide hints about their formation pathways. Recall that, based on previous observations (Dalessandro et al., 2013; Dattatrey et al., 2023), hotter (bluer), but relatively faint BSSs are likely to be formed through collisions, whereas slightly cooler (redder), but slightly brighter BSSs are likely to be formed by MT. Even though this is almost certainly not strictly followed by the universe (modulo stellar-mass loss rate, binary parameters, environmental factors, such as tidal shocks and disruption of binaries, etc.), it could be considered a first scenario for the formation of particular BSS sub-populations. Trying to provide a quantity that could give a hint about which pathway dominate certain evolutionary phases, JS21 defined a parameter called the “fractional mass excess”, or . This quantity is defined as
@@
To make headway in the discussion of BSS formation pathways, we study the behavior of as a function of cluster age and the equivalent BSS age from isochrone fitting, both of which are displayed in Figure 5 along with the BSS mass and mass offset scalings. We focus first on the left-side panels. In the top panel, we observe that, as expected, younger clusters host BSSs with higher masses, which decrease with cluster age. The Pearson coefficient also provide a strong correlation value of , combining both OCs and GCs. This correlation is consistent with the decreasing MSTO mass as a stellar population ages, since the lifetime of a star, , scales as , i.e. more massive stars have shorter lifetimes compared to less massive stars, which translates into the BSS mass relation in the top left panel of Figure 5. This correlation also appears stronger when only OCs are considered (), but this is not true for GC sample (). The reasons for GCs are the same as discussed in the previous sections.
@@
The middle panel on the left side of Figure 5 shows that the difference between the BSS mass and the MSTO mass as a function of cluster age. We see that the BSS with higher mass difference are absent towards the older cluster age with Pearson coefficient of . This is also expected from regular stellar evolution since more massive stars end their life disproportionately earlier while the MSTO mass decreases more slowly with increasing age, which is exactly what is observed, i.e. shrinks as time advances. We also notice here that GC sample alone does not show any correlation ().
@@
More surprisingly, we find no trend of with cluster-age (bottom left panel in Fig. 5) either for OCs or GCs with Pearson coefficient values and , respectively. However, we observe that the fraction of high- () BSSs is larger in OCs () compared to GCs (), which may indicate that the collisional formation channel is more pronounced in younger star clusters. This is turn might be related to sufficiently old dynamical ages of star cluster, which can, in principle, have undergone one or more core collapse, boosting the stellar encounter rates and, therefore, the formation of collision-induced BSS formation. However, this might be accomplished for star clusters at a few Gyr of age, and requires further investigation and modeling. We find and as the BSS fractions with low and high– (split at ), respectively, for OCs. In GCs, we find these proportions to be and for low and high–, respectively. Considering all clusters, the percentage of low– is and for high– we obtain .
@@
Additionally, we plot the BSS mass parameters vs. BSS ages in the right-hand side panels of Figure 5. The relations for and are consistent with those found against the cluster age (left side of the same Figure), i.e. both of these parameters decrease as a function of BSS age and stellar population age. In addition, we see various sequences that indicate that the BSSs in OCs and GCs predominantly populate different parameter spaces. A striking distribution is found in the against plot (bottom right panel) which shows two dominant sequences, both of which are consistent with decreasing as BSSs become older, albeit at different rates. We will later show that these sequences are likely representative of the different BSS formation channels. The pearson coefficients for all these three relations for OCs (GCs) are: () for vs BSS ages; () for vs BSS ages; and () for vs BSS ages.
@@
Figure 6 shows the percentage of low– and high– BSSs as a function of BSS age. For OCs, it is clear that the high– fraction vanishes above BSS ages of , whereas the low– fraction dominates, with a percentage in the same older age range. Curiously, the high– fraction in GCs increases up to for BSS ages between –, while the low– fraction experiences a corresponding dip. This may be due to the predominance of BSS formed through the collisional channel at those ages. Ferraro et al. (2009); Dalessandro et al. (2013); Simunovic et al. (2014) fitted collisional isochrones (Sills et al., 2009) to the blue sequence of BSSs, finding that their ages are , concluding that double sequences of BSSs in GCs may be products of violent interaction episodes in the past core-collapse events, which is in agreement with our age range found for BSS in GCs where the high– fraction increases and the low– fraction dips. For GCs older than the high– fraction rapidly vanishes similar to those for OCs. Overall, we observe for all BSSs in our sample low/high– total fraction is dominated by OCs for younger BSSs ages () and by GCs for older BSSs ages () as expected by their sample characteristics (see Sect. 3.1).

### v2

@@ 3.1 Absolute numbers and fractions of BSSs
Figure 3 shows the absolute number () of detected BSSs and their number ratio relative to MS stars () in each of our sample GCs and OCs as a function of the parent cluster age. For the denominator of , we use the number of MS stars in the luminosity range between the MSTO and MSTO+1 mag, similar to previous studies (JS21, Cordoni et al. 2023). We select MS stars using the same criteria for all clusters shown in the bottom panel of Figure 3 to guarantee a homogeneous comparison. Due to the larger average distances of GCs and therefore, their fainter MSTOs, our earlier photometric depth cut (see Sect. 2.2) collides with the fainter bound of the luminosity range definition for some distant GCs. For this reason we inspected by eye every CMD in our GC sample and only provide values for those clusters whose MS stars at fainter than the MSTO are within the photometric depth bounds. With this limitation, we are able to obtain values for 21 of 41 sample GCs, which is the reason why the number of GCs in the bottom panel of Figure 3 is smaller than the corresponding number in the upper panel. We also list the values of , and in Table 2 along with other stellar parameters of the clusters.
@@
The inset panels show zoom-ins at the oldest ages to better visualize any localized trends for the GC subsample. In Figure 3, we also compare the number of BSSs identified in this work against other Gaia-based studies from the literature, i.e. R21, JS21, Li et al. (2023) and Cordoni et al. (2023). We point out that our sample is the first that homogeneously studies BSSs in GCs and OCs. We observe that in all studies, including ours, the BSS population in OCs tends to increase in absolute and fractional numbers towards older cluster ages. This trend appears to reach a maximum at GC ages (), however, and it might even be consistent with the numbers of identified BSSs decreasing with increasing age slightly. We also perform sliding histograms of the distributions of () and (dashed line) in Figure 3. shows a clear trend of increasing with age, while the distribution appeared as a plateau in the range before suddenly rising again with the varying BSS fraction at older ages. We also perform the Pearson’s r-coefficients in the Monte Carlo method 222 Pearson’s test measures whether a pair of two variables are linearly correlated or not. The r-coefficient is defined as the covariance of the variables divided by the product of their standard deviations, and the value ranges from 1 to 1. 1 signifies perfect positive linear correlation between the variables, 1 signifies perfect linear anti-correlation, and 0 means no linear correlation. Based on Cohen (1988), coefficient values 0.5 (0.5) indicate a strong positive (negative) linear relation, values between 0.3 and 0.5 (0.3 and 0.5) indicate a moderate positive (negative) linear relation, and values between 0 and 0.3 (0 and 0.3) indicate a weak positive (negative) linear relation. The Pearson’s test also provides a p-value that indicates whether the observed correlation is statistically significant or not. It depends on the number of data points or the sample size. A smaller p-value (generally 0.05) suggests that the correlation is statistically significant, while for larger values (0.05), we can not conclude a significant correlation (Wasserstein & Lazar, 2016). (to include observational errors) for both distributions (Curran, 2014; Privon et al., 2020)333https://github.com/privong/pymccorrelation?tab=readme-ov-file and find that the values are 0.67 for and 0.45 for .444The numbers in all cases denote the correlation coefficient 50th percentile. The lower and upper errors denote the 16th and 75th percentiles, respectively. The p-values for both cases are found to be 0. The coefficients and p-values suggest that and , both correlate with the age of the clusters. Although the correlation for is moderate, one cannot deny the sudden rise of BSS fraction at . The plateau in in the range could be because, at a young age, massive stars reach larger radii in the AGB phase and mass transfer is possible even in wide binaries, which favours BSS formation to make a constant BSS fraction with respect to the MS population (Leiner & Geller, 2021). Other possible mechanisms of more BSS formation are wind accretion from the AGB companion and stable mass transfer in the subgiant branch or Hertzprung gap in younger clusters. The increase in the BSS fractions at older ages might be due to the longer MS lifetime of low-mass BSSs in older clusters and/or the contribution from collisional BSSs, however (Leiner & Geller, 2021, see their Fig.5).
@@
Pearson’s r-coefficient () suggests, however, that the anti-correlation is rather weak. A similar weak anti-correlation () is also found for as well. The p-values are found to be 0 for both cases. As the very central regions of the GCs are excluded in this study because of Gaia’s limitation, incorporating HST observations alongside Gaia data would be beneficial for investigating this correlation. At the same time, we need to include more number of GCs with deep photometry, as we missed to include some GCs due to the magnitude limitation of Gaia.
@@
3.2 Astrophysical parameters of BSSs
We derive the following astrophysical parameters for our sample BSSs: effective temperature (), stellar mass (), luminosity () and surface gravity (). We used three different approaches: the infrared flux method (IRFM; from color-temperature relations), the isochrone best–fit method (from PARSEC isochrones) and the XPP from Gaia DR3 spectra, if available. While the isochrone method provides all astrophysical parameters, the IRFM method provides and the XPP method gives & . We mainly use the parameters obtained from the isochrone method, but it is necessary to compare values, obtained from all three methods to check the consistency in the parameter estimations. In Appendix B, we describe all three methods and the comparison of the values obtained from them, in detail. In the following subsections, we summarize the results and compare the stellar parameters with the literature values.
@@ 3.2.1 Effective temperature
We notice that values extracted from different approaches range from 6000 K to 9000 K and appear consistent with each other with a little systematic offset towards higher temperature ( ). From these comparisons, we estimate that the systematic uncertainty between various methods is of the order K. From the IRFM we find for GCs and for OCs. From isochrone fitting we find for GCs and for OCs. Only 537 of 4399 BSSs ( of our sample) have XPP parameters measured for 279 stars in GCs and 258 stars in OCs. With this method, we find the mean BSS effective temperature of for GCs and for OCs. We list these parameters in Table 3, and the full table is available at the CDS.
Table 1 compare our derived values for with those found in the literature for various OCs and GC, as well as BSS samples. We find that our effective temperatures derived are in very good agreement with previous studies. For example, Jadhav et al. (2019), Sindhu et al. (2019) and Pandey et al. (2021) have analyzed different BSSs in the OC NGC 2682 (also known as M67) by fitting models to spectral energy distributions (SEDs) searching for BSS hot companions; these studies derive BSS temperatures in the range between and , which are in excellent agreement with our measurements, yielding a mean for all methods combined. We also point out the excellent agreement with the values derived in Gosnell et al. (2015), where the authors find for 19 BSSs, which is in very good agreement with our results independent of the applied method ( and , for the IRFM and isochrone method, respectively). Finally, our is also in very good agreement with theoretical predictions for BSSs in GCs from Stȩpień & Kiraga (2015) who calculate , and who had applied different initial conditions to binary systems.
@@ 3.2.2 Stellar mass and age of the BSSs
Stellar mass and age are important parameters in the study of the evolutionary state of BSSs. We derive these parameters along with equivalent stellar luminosity and stellar surface gravity for all BSSs in our sample based on the isochrone fitting method. We obtain the mean luminosity for GCs and for OCs, and find the mean surface gravity of for GCs and for OCs. These values correspond to a mean stellar mass for BSSs in GCs and for BSSs in OCs, which altogether translates to a mean stellar age of for BSSs in GCs and for BSSs in OCs.
We compare these values with previous results from the literature in Table 1. We find that the average mass of our BSS sample in GCs () differs from previous works, which report (Shara et al., 1997) and (Baldwin et al., 2016). This may be partly explained by the lower BSS numbers of previous studies, but also by either of the previous arguments, i.e. i) limitations of the Gaia data selection and ii) BSS mass segregation in GCs. In particular, the latter argument lends credence to the reason for this offset as the two previous studies are based on HST observations that target the GC centers. Our sample, on the other hand, is based on BSS samples from outside the half-light radius. Nevertheless, as can be appreciated in Table 1, our BSS mass results are in very good agreement with those measured by Simunovic (2017, ), Fiorentino et al. (2014, ) and Stȩpień & Kiraga (2015, ). The comparison with Simunovic (2017) is in particular striking, as this work measured the parameters for 1507 BSSs in 38 GCs using the isochrone fitting technique similar to our procedure, which makes it statistically the most robust comparison.
@@
Notes. Columns are as follows: 1) Cluster name. 2) Number of available BSSs in the literature with available astrophysical parameters. 3) Mean of the effective BSS temperature provided. 4) Mean of the decadal logarithm of the BSS surface gravity. 5) Mean of the BSS stellar mass. 6) Reference. For data provided by Jadhav & Subramaniam (2021, or JS21), we considered stars labeled BSS and pBSS (probable BSS).
@@
The average BSS mass for our OC sample ( compares favorably with the corresponding value found by JS21 considering the uncertainties. They report an average mass value of if only confirmed BSSs from their catalog are considered, and if both confirmed and candidate/probable BSSs listed in their catalog are included. As was explained previously, the difference between these works is likely the result of our Gaia magnitude cut-off criterion, discarding all bright stars with as they are likely saturated and therefore, their Gaia parameters likely affected by systematic errors. This not only explains the difference between our results, as well as why our dispersion in BSS mass for OCs is smaller, but also the lack of BSSs younger than Myr in our sample, most of which would naturally be among the most massive BSSs.
3.3 Effective temperature versus cluster age
@@
In Figure 4 we compare the average MSTO offset in of BSS in a cluster against the corresponding cluster age. To compare self-consistently all the BSSs with respect to the same evolutionary reference point, given the clusters’ very different ages and metallicities, we use the of their MSTO as a reference. For each BSS, we obtain . We then compute which is the mean difference between the BSS temperatures and the in each cluster. As XPP parameters are not available for stars at the MSTO, we only do this for obtained by isochrones and color–temperature relation, i.e. IRFM, and recall the systematic offsets found in the comparison of these methods (see Fig. 15). The error bars in represent the dispersion of BSS temperatures in each cluster, while the error bars on the horizontal axis are age uncertainties.
@@
In the top panel of figure 4, BSSs in OCs show on average a trend against the parent cluster age where younger OCs show larger offsets, i.e. hotter BSSs, with respect to those found in older OCs, or in other words, the difference between the effective temperature of BSSs and the MSTO effective temperature decreases with increasing parent cluster age. Pearson’s coefficients for IRFM (isochrone) method is (), suggesting the correlation to be weak, which could be due to a large dispersion in the BSS temperatures in the younger cluster. We expect the observed trend, however, because younger clusters host more massive stars which are also larger in diameter (e.g. Iben, 2013), and therefore, have likely more efficient MT between binary members and higher collision/capture probabilities with other stars. These hotter BSS stars evolve faster than BSS stars around the MSTO and, thus, the difference in average between BSSs and BSSs around MSTO is expected to decrease with time. We also notice that the dispersion in decreases with the age of the OC, which may indicate that BSSs tend to be more concentrated in the color-luminosity space as the OC ages, which is also shown in the bottom panel of the figure. Pearson’s coefficient values further confirm the trend: and for IRFM and isochrone methods, respectively.
@@
We do not observe any of these correlations for BSSs in GCs, which is also supported by the Pearson’s coefficient for the IRFM (isochrone) method () (see inset panels in Fig. 4). We recall, however, that we might be exposed to biases for the GC subsample: i) we applied a color cut when selecting BSSs in GCs to avoid including BHB stars (see Fig. 2 and Sect. 2.2) which, although unlikely may cut into the distribution of BSSs with extremely large (this could in principle be tested with UV photometry); ii) Gaia is only capable to observe stars outside (half-light radius) in GCs, which could change the mean if we assume that mass segregation has already acted on the BSS population (Contreras Ramos et al., 2012; Ferraro et al., 2012). Without consideration of these potential biases and taking the distributions in Figure 4 at face value, this means that as GCs age past Gyr, BSSs are more or less equally distributed in .
@@
In the middle panel of figure 4, we check if the fractional excess with respect to has any correlation with the cluster age. We do not find any trend for either OCs or GCs, which is also supported by the Pearson’s coefficient test. Pearson’s coefficients for IRFM (isochrone) are () for OCs and ().
3.4 Fractional mass excess of BSSs
@@
BSSs are expected to gain mass during their MS lifetime. This means that the “excess” between their current mass and the stellar mass at the MSTO could provide hints about their formation pathways. Recall that, based on previous observations (Dalessandro et al., 2013; Dattatrey et al., 2023), hotter (bluer), but relatively faint BSSs are likely to be formed through collisions, whereas slightly cooler (redder), but slightly brighter BSSs are likely to be formed by MT. Even though this is almost certainly not strictly followed by the universe (modulo stellar-mass loss rate, binary parameters, environmental factors, such as tidal shocks and disruption of binaries, etc.), it could be considered a first scenario for the formation of particular BSS subpopulations. Trying to provide a quantity that could give a hint which pathway dominates certain evolutionary phases, JS21 defined a parameter called the “fractional mass excess”, or . This quantity is defined as
@@
To make headway in the discussion of BSS formation pathways, we study the behavior of as a function of cluster age and the equivalent BSS age from isochrone fitting, both of which are displayed in Figure 5 along with the BSS mass and mass offset scalings. We focus first on the left-side panel. In the top panel, we observe that, as expected, younger clusters host BSSs with higher masses, which decrease with cluster age. The Pearson coefficient also provides a strong correlation value of , combining both OCs and GCs. This correlation is consistent with the decreasing MSTO mass as a stellar population ages, since the lifetime of a star, , scales as , i.e. more massive stars have shorter lifetimes compared to less massive stars, which translates into the BSS mass relation in the top left panel of Figure 5. This correlation also appears stronger when only OCs are considered (), but this is not true for the GC sample (). The reasons for GCs are the same as discussed in the previous sections.
@@
The middle panel on the left side of Figure 5 shows the difference between the BSS mass and the MSTO mass as a function of cluster age. We see that the BSS with higher mass difference are absent towards the older cluster age with Pearson coefficient of . This is also expected from regular stellar evolution since more massive stars end their life disproportionately earlier while the MSTO mass decreases more slowly with increasing age, which is exactly what is observed, i.e. shrinks as time advances. We also notice here that GC sample alone does not show any correlation ().
@@
More surprisingly, we find no trend of with cluster-age (bottom left panel in Fig. 5) either for OCs or GCs with Pearson coefficient values and , respectively. We observe, however, that the fraction of high- () BSSs is larger in OCs () compared to GCs (), which may indicate that the collisional formation channel is more pronounced in younger star clusters. This in turn might be related to sufficiently old dynamical ages of star cluster, which can, in principle, have undergone one or more core collapse, boosting the stellar encounter rates, and therefore, the formation of collision-induced BSS formation. This might be accomplished for star clusters with ages of a few billion years, however, and it requires further investigation and modeling. We find and as the BSS fractions with low and high– (split at ), respectively, for OCs. In GCs, we find these proportions to be and for low and high–, respectively. Considering all clusters, the percentage of low– is and for high– we obtain .
@@
Additionally, we plot the BSS mass parameters versus BSS ages in the right panels of Figure 5. The relations for and are consistent with those found against the cluster age (left side of the same Figure), i.e., both of these parameters decrease as a function of BSS age and stellar population age. In addition, we see various sequences that indicate that the BSSs in OCs and GCs predominantly populate different parameter spaces. A striking distribution is found in the against plot (bottom right panel) which shows two dominant sequences, both of which are consistent with decreasing as BSSs become older, albeit at different rates. We will later show that these sequences are likely representative of the different BSS formation channels. The pearson coefficients for all these three relations for OCs (GCs) are: () for versus BSS ages; () for versus BSS ages; and () for versus BSS ages.
@@
Figure 6 shows the percentage of low– and high– BSSs as a function of BSS age. For OCs, it is clear that the high– fraction vanishes above the BSS ages of , whereas the low– fraction dominates, with a percentage in the same older age range. Curiously, the high– fraction in GCs increases to for BSS ages between –, while the low– fraction experiences a corresponding dip. This may be due to the predominance of BSS formed through the collisional channel at those ages. Ferraro et al. (2009); Dalessandro et al. (2013); Simunovic et al. (2014) fitted collisional isochrones (Sills et al., 2009) to the blue sequence of BSSs, finding that their ages are , concluding that double sequences of BSSs in GCs may be products of violent interaction episodes in the past core-collapse events, which is in agreement with our age range found for BSS in GCs where the high– fraction increases and the low– fraction dips. For GCs older than the high– fraction rapidly vanishes similar to those for OCs. Overall, we observe that for all BSSs in our sample low/high– total fraction is dominated by OCs for younger BSSs ages () and by GCs for older BSSs ages () as expected by their sample characteristics (see Sect. 3.1).